Identify the solutions of the systems of equations, if any
-3x-4y=2
8y=-6x-4
Thanks

Answer by josgarithmetic(39618)

Put the second equation, 8y=-6x-4, into the same form as the first and compare:
8y+6x=-4
4y+3x=-2
3x+4y=-2

Note how the coefficents on x and y are in proportion to the ones in the first equation. NOW, try multiplying both sides of the second equation (now transformed) by negative 1:

-3x-4y=2.
This is exactly the same as the first equation. The original system is really one single equation written in two different ways. INFINITE SOLUTIONS.